# L-WISE: Boosting human visual category learning through model-based image selection and enhancement

### v2

@@ 1 Introduction
Figure 1: Robustified ANNs can be used out-of-the-box as image recognition difficulty estimators and ground truth percept enhancers. We consider a 16-way basic animal classification task. Panel A1 shows the correspondence between human categorization accuracy and model-computed ground truth logit activation values. The curve denotes a logistic regression model predicting the probability of a correct response using only the logit value ($p<0.001$ from the Wald statistic, $\text{AUC}=0.72$ under 10-fold cross validation). A2 shows example images with varying ground truth logit values (predicted difficulty). B1 shows how perturbing images via ground truth logit maximization increases human recognition accuracy progressively with the $\ell_{2}$-norm perturbation pixel budget $\epsilon$. Other off-the-shelf image enhancement methods do not increase categorization accuracy, despite inducing larger perturbations of $\epsilon=43$, $\epsilon=106$, and $\epsilon=26$ on average from left to right. B2 shows example images: unmodified (left), enhanced by ground truth logit maximization with pixel budgets ${\epsilon}=10$ and ${\epsilon}=20$ (middle), and enhanced by baseline off-the-shelf methods (right). All vertical error bars are 95% confidence intervals by bootstrap. Horizontal error bars (panel A1) show the standard deviation among images within each logit value bin.
@@
Beyond categorization of familiar visual categories, one practically important task is learning to recognize new, unfamiliar categories. Although humans can readily learn many new categories even from a single example (Lake et al., 2011), some consequential tasks require extensive training to reach high levels of performance. For example, medical specialists such as pathologists and radiologists devote numerous hours to mastering the diagnosis of various diseases from medical images. We identify a potential strategy to accelerate the visual learning process by extrapolating from the perceptual learning literature. Simple visual tasks, such as line orientation discrimination, reveal a curriculum effect whereby providing easy examples to a novice human learner before gradually increasing the difficulty promotes faster perceptual learning (Lu & Dosher, 2022). Motivated by these findings, we ask whether the human-aligned nature of robustified ANNs allows them to augment human learning in complex image domains, chiefly by reducing the initial task difficulty and then increasing the difficulty as learning progresses. A demonstration that these models can be used to enhance learning serves as an additional scientific test of the models’ alignment with human perception, while also pioneering a potentially beneficial application of ANNs in education.
@@
To test the viability of enhancing image category learning in humans, we first establish two key empirical observations, summarized in Fig. 1: (i) we found that the human error rate in an image categorization task is strongly predicted by the ground truth logit activation of a robustified ANN, making it a valid image recognition difficulty score for humans; (ii) we found that this relationship also holds in reverse – pixel-level perturbations can be generated by the model to maximize the ground truth logit activation, producing an “enhanced” version of the image that is easier for humans to recognize as the ground truth category. While previous findings demonstrate that model-guided perturbations can modulate human perception away from the ground truth category (Gaziv et al., 2023), we conversely seek to amplify category-relevant features to facilitate correct classification. We observe that our model-based image enhancements yield significant increases in human accuracy on a classification task with basic animal categories, unlike conventional image enhancement techniques that adjust low-level image properties such as lighting and color-contrast. We thus propose an algorithm that combines model-based image enhancement and image difficulty prediction to generate optimized curricula for novice humans learning challenging image categorization tasks.
@@
Our proposed method, “Logit-Weighted Image Selection and Enhancement” (L-WISE), is illustrated in Fig. 2. In our primary experimental setting, a human participant learns an unfamiliar image categorization task by viewing a series of examples, providing a category judgment on each image before receiving feedback indicating the correct category (Fig. 2A,B). Upon completion of the training phase, test accuracy is measured on held-out images in similar trials without feedback (Fig. 2C). L-WISE intervenes on this naive visual learning baseline by using a robustified ANN model in two ways: (i) it uses the ground truth logit difficulty score to sample training images based on a predefined, increasing schedule of maximal difficulty per trial (Fig. 2D,E); (ii) it enhances training images with a perturbation magnitude that decreases along a similarly predefined schedule (Fig. 2D,F).
@@ 2 Related Work
Our primary goal is to apply difficulty prediction and image enhancement to augment human learning. The emerging field of machine teaching (Zhu, 2015) employs machine learning to find or generate optimal “teaching sets” that can be used to train models or humans. While many such approaches have been successfully applied to training machine learning models (e.g., Liu et al. (2017); Qiu et al. (2023)), few studies have successfully enhanced image category learning in humans and most of these focus on teaching set selection. Singla et al. (2014) propose STRICT, which optimizes the expected decrease in learner error based on how the selected images and their labels constrain a linear hypothesis class in a feature space. Johns et al. (2015) extend a similar approach to select images in an online fashion by modeling the learner’s progress. MaxGrad (Wang et al., 2021a) uses bi-level optimization to iteratively refine a teaching set by modeling learners as optimal empirical risk minimizers. Most similar to our work are approaches like EXPLAIN (Mac Aodha et al., 2018), which uses ANN class activation maps (CAMs) to highlight relevant image regions while providing feedback to the learner. EXPLAIN also selects a curriculum of images based on (a) a multi-class adaptation of STRICT, (b) representativeness (mean feature-space distance to other images of the same class), and (c) the estimated difficulty (entropy) of the CAM explanations. Chang et al. (2023) use bounding boxes to highlight image regions attended to by experts and not novices, allowing humans to more accurately match bird or flower images to one species among five shown in a gallery.
Our approach to learning augmentation departs from previous studies in several ways. We make explicit estimates of image difficulty with unprecedented accuracy to select easier images for early-stage learners. Our work is unique in employing category-specific image enhancement, which is a novel technique in itself, to improve the teaching efficacy of a given set of images. While Mac Aodha et al. (2018) and Chang et al. (2023) help learners by explicitly highlighting where learners should attend to in each image, we take a distinct and complementary approach by implicitly highlighting what learners must attend to in order to classify images correctly.
@@ 3 Overview of approach and experiments
Our approach to improving visual learning in humans is based on two key observations regarding human-aligned ANN models of the ventral visual stream: (i) they can accurately predict the recognition difficulty of specific images for humans (Fig. 1A), and (ii) they can be used to perturb images in a way that enhances the ability of humans to accurately report the original ground truth category (Fig. 1B). In other words, these models can be used out-of-the-box as category-recognition difficulty estimators, and category-percept enhancers. As such, we propose using them to design sequences of images that humans can use to more efficiently learn to recognize unfamiliar image categories. We test our approach in a variety of challenging image domains: natural photographs (moth species classification), dermoscopy images (skin lesion classification), and histology images (benign vs. pre-cancerous classification).
@@ 3.1 Training task-specific robustified models
To obtain robustified models for a task specific category space, we adversarially trained ResNet-50 ANNs (He et al., 2016) on the ImageNet-1K (Deng et al., 2009) and iNaturalist 2021 (Van Horn et al., 2021) datasets (separately) using the same techniques as Mądry et al. (2018). To adapt the resulting model to the three categorization tasks of interest, we conducted additional adversarial fine-tuning on the corresponding smaller datasets: a subset of moth species images from iNaturalist (from iNaturalist pretraining), the HAM10000 dermoscopy dataset (Tschandl et al., 2018) (ImageNet pretraining), and the MHIST histology dataset (Wei et al., 2021) (ImageNet pretraining).
@@ 3.4 Boosting image category learning in humans
Harnessing key observations of robustified ANNs, our “L-WISE” approach uses one such model to optimize the training phase in a way that improves the test accuracy via two mechanisms: (i) sampling training images based on their predicted recognition difficulty, and (ii) enhancing the training images. We adjust the “strength” of both mechanisms in a time-dependent manner: for the former, we set the maximum allowable difficulty score for an image to be shown in a given trial, and for the latter, we limit the enhancement perturbation magnitude to an $\ell_{2}$ norm pixel-budget $\epsilon$. The user of our approach can flexibly define arbitrary time-dependent profiles for image selection and enhancement (Fig. 2D-F). In this study, we focused on a linear ramp schedule for the allowable image difficulty at a given time, and on exponential tapering of the enhancement $\epsilon$. Intuitively, this should correspond to an easy-to-challenging trend during the training phase. Extensively optimizing these schedules was not a goal of our study, which rather serves as a proof-of-concept. Incorporating image selection and enhancement led to significant gains in the accuracy of human participants on unmodified, randomly-selected test images, while also reducing the time needed to complete the training phase (which consists of a constant number of trials). This result was robustly observable across the varied image domains and category spaces we tested.
@@ 4 Results
Figure 3: Novice learners who had their curriculum augmented by our method showed improved test-time categorization accuracy for previously unfamiliar categories. This figure shows empirical results for a 4-way fine-grained moth species classification task. Panel A shows examples of the four moth classes, side-by-side with their model-enhanced versions at the highest pixel budget used in our experiments ($\epsilon=8$). While subtle, one notable difference is the distinctive wing spots of moth class 2, which are enlarged in the enhanced version of the image. Also included are difference images showing the (5x magnified) difference between original and enhanced versions, and heat maps with more red coloration in regions of larger changes from enhancement. B compares the average smoothed accuracy of participants in the L-WISE group and a control group. Shaded areas denote 95% CI by bootstrap. The test accuracy gain of the L-WISE group relative to the control group is statistically significant ($\chi^{2}(1)$ test, $p<0.001$). C,D show the trial-dependent empirical profiles of the average image difficulty percentile of selected images, which (noisily) increases step-wise, and the perturbation pixel budget for enhancement ($\epsilon$), which decreases step-wise. The profiles of the baseline controls (black dotted lines) are uniform, indicating randomly-selected images with no enhancement.
@@
We used robustified ANNs to both enhance images and predict the difficulty of images across multiple domains. We applied both of these techniques to improve the final test performance (on original, unmodified images) of novices learning challenging image classification tasks.
@@ 4.1 Robust models can both predict image recognition difficulty and reduce it
We tested the effects of a novel model-based image enhancement algorithm on human image categorization accuracy. We demonstrate that we can enhance images by maximizing the ground truth logit from a robustified ANN (ResNet-50) using gradient ascent in image pixel space. As the magnitude of the enhancement perturbations grows ($\ell_{2}$ norm pixel budget $\epsilon$), human participants become increasingly accurate on a 16-way animal photograph classification task derived from ImageNet (Fig. 1B, chance = 1/16). While mean accuracy on the original, unmodified ($\epsilon=0$) images was 0.75, mean accuracy on enhanced images was as high as 0.84 at $\epsilon=20$. The accuracy gains from enhancement appear to approach a saturation point as the perturbations grow larger. The improvements in accuracy are also somewhat dependent on the starting ground truth logit score, as shown in Appendix Fig. S11: accuracy gains are larger for “difficult” images than for “easy” images. Baseline enhancement algorithms Contrast-Limited Adaptive Histogram Equalization (CLAHE, Zuiderveld (1994)), Multi-Scale Retinex with Color Restoration (MSRCR, Jobson et al. (1997); Petro et al. (2014)), and Adobe Photoshop Lightroom’s “Auto” enhancement feature (LR, Adobe Inc. (2024)) had no significant effect on human accuracy, despite inducing image perturbations of considerably larger $\ell^{2}$ norm on average than the $\ell^{2}$ pixel budget $\epsilon$ values we used for model-based enhancement.
@@
4.2 L-WISE improves both learning speed and test accuracy for humans
We applied both image difficulty prediction and image enhancement as part of a novel method that designs curriculum image sequences for novices learning challenging image classification tasks. Our proposed algorithm, “Logit-Weighted Image Selection and Enhancement” (L-WISE), operates on image trial sequences used to train human participants on image classification tasks through trial-by-trial feedback. The performance of each participant is evaluated in a subsequent testing phase without feedback, which includes only randomly-selected, unmodified images (unaffected by L-WISE). During the early portion of the training phase, L-WISE randomly selects images from below a certain difficulty percentile that linearly increases as the training phase progresses. Selected images are enhanced at each trial during this period, within an $\ell_{2}$ pixel budget $\epsilon$ that decreases in a stepwise-exponential fashion (Figs. 2E-F and 3C-D).
@@ 4.3 Image enhancement and selection both contribute to efficacy of L-WISE
Idaea moth photos Skin lesion dermoscopy
@@
Table 1: Both image enhancement tapering (ET) and image difficulty selection (DS) contribute to the ability of L-WISE to assist learners. The benefits of image enhancement are dependent on easy-to-hard sequencing (“ET” outperforms “ET (shuffled)”), but the benefits of difficulty-based selection appear to stem from simply showing an easier distribution of images during training (“DS (shuffled)” performs as well as or better than “DS”). Training durations are in minutes. Parentheses show 95% bootstrap confidence intervals for the mean with 10,000 bootstrap replicates. * denotes a significant difference in accuracy from the control group ($p<0.01$, $\chi^{2}(1)$ test).
We tested several ablated versions of L-WISE to determine the relative contributions of its components, which include (A) image enhancement based on logit maximization, (B) selection of relatively easy images according to logit-estimated difficulty, and (3) easy-to-hard curriculum trends enabled by A and B. The results of our ablation study, which employed the moth classification and skin lesion dermoscopy tasks, are shown in Table 1. In “Enhancement Tapering,” (ET) only the enhancement component of L-WISE is used, with no difficulty-based image selection. Conversely, in “Difficulty Selection” (DS), images are selected based on difficulty but not enhanced. In “ET (shuffled)” and “DS (shuffled),” the order of training trials is randomly permuted after the ET or DS intervention respectively. We hypothesized that shuffling would abolish at least some of the potential accuracy gains attributable to ET and DS due to the possible helpfulness of easy-to-hard trends.

### v4

@@ 1 Introduction
Figure 1: Robustified ANNs can be used out-of-the-box as image recognition difficulty estimators and ground truth percept enhancers. We consider a 16-way basic animal classification task. Panel A1 shows the correspondence between human categorization accuracy and model-computed ground truth logit activation values. The curve denotes a logistic regression model predicting the probability of a correct response using only the logit value ($p<0.001$ from the Wald statistic, $\text{AUC}=0.72$ under 10-fold cross validation). A2 shows example images with varying ground truth logit values (predicted difficulty). B1 shows how perturbing images via ground truth logit maximization increases human recognition accuracy progressively with the $\ell_{2}$-norm perturbation pixel budget $\epsilon$. Other off-the-shelf image enhancement methods do not increase categorization accuracy, despite inducing larger perturbations of $\epsilon=43$, $\epsilon=106$, and $\epsilon=26$ on average from left to right. B2 shows example images: unmodified (left), enhanced by ground truth logit maximization with pixel budgets ${\epsilon}=10$ and ${\epsilon}=20$ (middle), and enhanced by baseline off-the-shelf methods (to the right of the dotted line). All vertical error bars are 95% confidence intervals by bootstrap. Horizontal error bars in panel A1 show the standard deviation among images within each logit value bin.
@@
Beyond categorization of familiar visual categories, one practically important task is learning to recognize new, unfamiliar categories. Although humans can readily learn many new categories even from a single example (Lake et al., 2011), some consequential tasks require extensive training to reach high levels of performance. For example, medical specialists such as pathologists and radiologists devote numerous hours to mastering the diagnosis of various diseases from medical images. We identify a potential strategy to accelerate the visual learning process by extrapolating from the perceptual learning literature. Simple visual tasks, such as line orientation discrimination, reveal a curriculum effect whereby providing easy examples to a novice human learner, before gradually increasing the difficulty, promotes faster perceptual learning (Lu & Dosher, 2022). Motivated by these findings, we ask whether the human-aligned nature of robustified ANNs allows them to augment human learning in complex image domains, chiefly by reducing the initial task difficulty and then increasing the difficulty as learning progresses. A demonstration that these models can be used to enhance learning serves as an additional scientific test of the models’ alignment with human perception, while also pioneering a potentially beneficial application of ANNs in education.
@@
To test the viability of enhancing image category learning in humans, we first establish two key empirical observations, summarized in Fig. 1: (i) we find that the human error rate in an image categorization task is strongly predicted by the ground truth logit activation value of a robustified ANN, making it a valid image recognition difficulty score for humans; (ii) we find that this relationship also holds in reverse – pixel-level perturbations can be generated by the model to maximize the ground truth logit activation, producing an “enhanced” version of the image that is easier for humans to recognize as the ground truth category. While previous findings demonstrate that model-guided perturbations can modulate human perception away from the ground truth category (Gaziv et al., 2023), we conversely seek to amplify category-relevant features to facilitate correct classification. We observe that our model-based image enhancements yield significant increases in human accuracy on a classification task with basic animal categories, unlike conventional image enhancement techniques that adjust low-level image properties such as lighting and contrast. We thus propose an algorithm that combines model-based image enhancement and image difficulty prediction to generate optimized curricula for novice humans learning challenging image categorization tasks.
@@
Our proposed method, “Logit-Weighted Image Selection and Enhancement” (L-WISE), is illustrated in Fig. 2. In our primary experimental setting, a human participant learns an unfamiliar image categorization task by viewing a series of examples, providing a category judgment for each image before receiving feedback indicating the correct category (Fig. 2A,B). Upon completion of the training phase, test accuracy is measured on held-out images in similar trials without feedback (Fig. 2C). L-WISE intervenes on this naive visual learning baseline by using a robustified ANN model in two ways: (i) it uses the ground truth logit difficulty score to sample training images based on a predefined, increasing schedule of maximal difficulty per trial (Fig. 2D,E); (ii) it enhances training images with a perturbation magnitude that decreases along a similarly predefined schedule (Fig. 2D,F).
@@ 2 Related Work
Our primary goal is to apply difficulty prediction and image enhancement to augment human learning. The emerging field of machine teaching (Zhu, 2015) employs machine learning to find or generate optimal “teaching sets” that can be used to train models or humans. While many such approaches have been successfully applied to training machine learning models (e.g., Liu et al. (2017); Qiu et al. (2023)), few studies have successfully enhanced image category learning in humans and most of these focus on teaching set selection. Singla et al. (2014) propose STRICT, which optimizes the expected decrease in learner error based on how the selected images and their labels constrain a linear hypothesis class in a feature space. Johns et al. (2015) extend a similar approach to select images in an online fashion by modeling the learner’s progress. MaxGrad (Wang et al., 2021a) uses bi-level optimization to iteratively refine a teaching set by modeling learners as optimal empirical risk minimizers. Most similar to our work are approaches like EXPLAIN (Mac Aodha et al., 2018), which uses ANN class activation maps (CAMs) to highlight relevant image regions while providing feedback to the learner. EXPLAIN also selects a curriculum of images based on (i) a multi-class adaptation of STRICT, (ii) representativeness (mean feature-space distance to other images of the same class), and (iii) the estimated difficulty (entropy) of the CAM explanations. Chang et al. (2023) use bounding boxes to highlight image regions attended to by experts and not novices, allowing humans to more accurately match bird or flower images to one species among five shown in a gallery.
Our approach to learning augmentation departs from previous studies in several ways. We make explicit estimates of image difficulty with unprecedented accuracy to select easier images for early-stage learners. Our work is unique in employing category-specific image enhancement, which is a novel technique in itself, to improve the teaching efficacy of a given set of images. While Mac Aodha et al. (2018) and Chang et al. (2023) help learners by explicitly highlighting where learners should attend to in each image, we take a distinct and complementary approach by implicitly highlighting what learners should attend to in order to classify images correctly.
@@ 3 Overview of approach and experiments
Our approach to improving visual learning in humans is based on two key observations regarding human-aligned ANN models of the ventral visual stream: (i) they can accurately predict the recognition difficulty of specific images for humans (Fig. 1A), and (ii) they can be used to perturb images in a way that enhances the ability of humans to accurately report the original ground truth category (Fig. 1B). In other words, these models can be used out-of-the-box as category-recognition difficulty estimators and category-percept enhancers. As such, we propose using them to design sequences of images that humans can use to more efficiently learn to recognize unfamiliar image categories. We test our approach in a variety of challenging image domains: natural photographs (moth species classification), dermoscopy images (skin lesion classification), and histology images (benign vs. pre-cancerous colon tissue classification).
@@ 3.1 Training task-specific robustified models
To obtain robustified models for task-specific category spaces, we adversarially trained ResNet-50 ANNs (He et al., 2016) on the ImageNet-1K (Deng et al., 2009) and iNaturalist 2021 (Van Horn et al., 2021) datasets (separately) using the same techniques as Ma̧dry et al. (2018). To adapt the resulting model to the three categorization tasks of interest, we conducted additional adversarial fine-tuning on the corresponding smaller datasets: a subset of moth species images from iNaturalist (after iNaturalist pretraining), the HAM10000 dermoscopy dataset (Tschandl et al., 2018) (ImageNet pretraining), and the MHIST histology dataset (Wei et al., 2021) (ImageNet pretraining).
@@ 3.4 Boosting image category learning in humans
Harnessing key observations of robustified ANNs, our “L-WISE” approach uses one such model to optimize the training phase in a way that improves the test accuracy via two mechanisms: (i) sampling training images based on their predicted recognition difficulty, and (ii) enhancing the training images. We adjust the “strength” of both mechanisms in a time-dependent manner: for the former, we set the maximum allowable difficulty score for an image to be shown in a given trial, and for the latter, we limit the enhancement perturbation magnitude to an $\ell_{2}$ norm pixel-budget $\epsilon$. The user of our approach can flexibly define arbitrary time-dependent profiles for image selection and enhancement (Fig. 2D-F). In this study, we used a step-wise linear ramp schedule for the allowable image difficulty at a given time, and exponential tapering of the enhancement $\epsilon$. Intuitively, this should correspond to an easy-to-challenging trend during the training phase. Extensively optimizing these schedules was not a goal of our study, which serves primarily as a proof-of-concept. Applying image selection and enhancement led to significant gains in the accuracy of human participants on unmodified, randomly-selected test images, while also reducing the time needed to complete the training phase (which consists of a constant number of trials). This result was robustly observable across the varied image domains and category spaces we tested.
@@ 4 Results
Figure 3: Novice learners who had their curriculum augmented by our method showed improved test-time categorization accuracy for previously unfamiliar categories. This figure shows empirical results from a 4-way fine-grained moth species classification task. Panel A shows examples of the 4 moth classes, side-by-side with their model-enhanced versions at the highest pixel budget used in our experiments ($\epsilon=8$). While subtle, one notable difference is the distinctive wing spots of moth class 2, which are enlarged in the enhanced version of the image. Also included are difference images showing the (5x magnified) difference between original and enhanced images, and heat maps with more red coloration in regions of larger changes from enhancement. B compares the average smoothed accuracy of participants in the L-WISE group and a control group. Shaded areas denote the standard error of the mean. The test accuracy gain of the L-WISE group relative to the control group is statistically significant ($\chi^{2}(1)$ test, $p<0.001$). C, D show the trial-dependent empirical profiles of the average image difficulty percentile of selected images, which (noisily) increases step-wise, and the perturbation pixel budget for enhancement ($\epsilon$), which decreases step-wise. These profiles are uniform in the control group (black dotted lines), denoting randomly-chosen non-enhanced images.
@@
We used robustified ANNs to both enhance images and predict the difficulty of images across multiple domains. We applied both of these techniques to improve the final test performance (on unmodified, randomly-chosen images) of novice humans learning challenging image classification tasks.
@@ 4.1 Robust models can both predict image recognition difficulty and reduce it
We tested the effects of a novel model-based image enhancement algorithm on human image categorization accuracy. We demonstrate that we can enhance images by maximizing the ground truth logit from a robustified ANN (ResNet-50) using gradient ascent in image pixel space. As the magnitude of the enhancement perturbations grows ($\ell_{2}$ norm pixel budget $\epsilon$), human participants become increasingly accurate on a 16-way animal photograph classification task derived from ImageNet (Fig. 1B, chance = 1/16). While mean accuracy on the original, unmodified ($\epsilon=0$) images was 0.75, mean accuracy on enhanced images was as high as 0.84 (at $\epsilon=20$). The accuracy gains from enhancement appear to approach a saturation point as the perturbations grow larger. The improvements in accuracy are also somewhat dependent on the starting ground truth logit score, as shown in Appendix Fig. S11: accuracy gains are larger for “difficult” images than for “easy” images. Baseline enhancement algorithms Contrast-Limited Adaptive Histogram Equalization (CLAHE, Zuiderveld (1994)), Multi-Scale Retinex with Color Restoration (MSRCR, Jobson et al. (1997); Petro et al. (2014)), and Adobe Photoshop Lightroom’s “Auto” enhancement feature (LR, Adobe Inc. (2024)) had no significant effect on human accuracy, despite inducing image perturbations of considerably larger $\ell_{2}$ norm on average than the $\ell_{2}$ pixel budget $\epsilon$ values we used for model-based enhancement.
@@
4.2 L-WISE improves both test accuracy and learning speed for humans
We applied both image difficulty prediction and image enhancement as part of a novel method that designs image sequence curricula for novices learning challenging image classification tasks. Our proposed algorithm, “Logit-Weighted Image Selection and Enhancement” (L-WISE), operates on image trial sequences used to train human participants on image classification tasks through trial-by-trial feedback. The performance of each participant is evaluated in a subsequent testing phase without feedback, which includes only randomly-selected, unmodified images (unaffected by L-WISE). During the early portion of the training phase, L-WISE randomly selects images from below a certain difficulty percentile that stepwise-linearly increases as the training phase progresses. Selected images are enhanced at each trial during this period, via perturbations within an $\ell_{2}$ pixel budget $\epsilon$ that decreases in a stepwise-exponential manner (Figs. 2E-F and 3C-D).
@@ 4.3 Image enhancement and selection both contribute to efficacy of L-WISE
Idaea moth photos Skin lesion dermoscopy images
@@
Table 1: Both image enhancement tapering (ET) and image difficulty selection (DS) contribute to the ability of L-WISE to assist learners. The benefits of image enhancement are dependent on easy-to-hard sequencing (“ET” outperforms “ET (shuffled)”), but the benefits of difficulty-based selection appear to stem from simply showing an easier distribution of images during training (“DS (shuffled)” performs as well as or better than “DS”). Training durations are in minutes. Values in parentheses show 95% confidence intervals from 10,000 bootstrap replicates. * denotes significant differences in accuracy from the control group ($p<0.01$, $\chi^{2}(1)$ test).
We tested several ablated versions of L-WISE (in the moth and dermoscopy tasks) to determine the relative contributions of its components: (A) image enhancement based on logit maximization, (B) selection of images according to logit-estimated difficulty, and (C) easy-to-hard curriculum trends enabled by A and B (Table 1). In “Enhancement Tapering” (ET), only the enhancement component of L-WISE is active, with random image selection as in the control group. Conversely, in “Difficulty Selection” (DS), images are selected based on difficulty but not enhanced. In “ET (shuffled)” and “DS (shuffled),” after applying ET or DS, the ordering of affected training trials is randomly permuted. Shuffling flattens easy-to-hard trends, isolating effects from (i) the mere presence of enhanced images in ET, and (ii) seeing easier images on average in DS (DS limits max. difficulty of early images, so DS/DS (shuffled) have easier training images than Control on average; see Fig. 3C).